OpenStudy (anonymous):

what would be the first step to get the derivative of y=1-cos2x+2cos^2x

zepdrix (zepdrix):

Mmm so what are you stuck on? The trig function got you confused?

zepdrix (zepdrix):

trig functions*

zepdrix (zepdrix):

We'll have to apply chain rule in each case.

OpenStudy (anonymous):

so both are 2 different chain rules so then it would be 1-2cos2x but i dont know what it would be for 2cos^2x because of it's exponent on cos

zepdrix (zepdrix):

Woops you applied the chain, the derivative of the inner function (2x). But you never took the derivative of the outer function, cosine.

zepdrix (zepdrix):

\[\Large\bf\sf \left(-\cos 2x\right)'\quad=\quad \sin2x\color{royalblue}{(2x)'}\quad=\quad 2 \sin 2x\]

zepdrix (zepdrix):

The derivative of a `constant` is zero. So the 1 differentiates to 0.

OpenStudy (anonymous):

ok i see now, but the 2nd part still confuses me

zepdrix (zepdrix):

It's confusing because of the power notation we use on trig functions. It makes it difficult to see what the `outermost` function truly is.\[\Large\bf\sf 2\cos^2x\quad=\quad 2(\cos x)^2\]From this form we can see that we need to apply the power rule to the outer function first.

zepdrix (zepdrix):

\[\Large\bf\sf \frac{d}{dx}2(\cos x)^2\quad=\quad 2\cdot 2(\cos x)^1\color{royalblue}{\frac{d}{dx}\cos x}\]

zepdrix (zepdrix):

Understand that first step in differentiating it? I applied the power rule to the 2 exponent. The blue part is due to the chain rule, we need to find the derivative of that still.

OpenStudy (anonymous):

Now I see, it would be best to just rewrite it like that, thanks so much!